# LRPO: Enhancing Blind Face Restoration through Online Reinforcement Learning

### v1

@@ 1 Introduction
Blind Face Restoration (BFR), which aims to reconstruct high-quality (HQ) faces from low-quality (LQ) inputs with unknown degradations, has made rapid progress in recent years. Modern BFR methods typically exploit various types of priors to establish direct mappings from LQ to HQ. These priors can be categorized into several types: (1) geometric priors (e.g., facial landmarks (Chen et al., 2018; Kim et al., 2019), parsing maps (Chen et al., 2021), and component heatmaps (Yu et al., 2018)) that provide structural guidance; (2) generative priors (Wang et al., 2021; Chan et al., 2021; Yang et al., 2021) from pre-trained models like StyleGAN (Karras et al., 2019; 2020) that enable realistic detail reconstruction; (3) discrete codebook priors (Gu et al., 2022; Zhou et al., 2022) improve restoration fidelity; and (4) diffusion priors (Wu et al., 2024; Lin et al., 2024; Chen et al., 2024; Yue & Loy, 2024; Wang et al., 2023b) that have recently attracted significant attention. Diffusion models offer distinct advantages including robust generative capability, stable optimization, and superior control over output diversity, making them particularly effective for producing high-quality, visually pleasing face restorations.
However, despite the advantages of diffusion priors, BFR remains fundamentally challenging. The task is inherently an ill-posed inverse problem where a single LQ input can correspond to multiple plausible HQ solutions, making it difficult to determine the optimal restoration. Current methods are constrained by their deterministic nature–they learn a fixed one-to-one mapping that produces a single output without considering alternative solutions. This lack of exploration within the vast solution space prevents these methods from discovering potentially superior restorations, leading to suboptimal results (Zhou et al., 2021).
To address these exploration limitations, we propose incorporating reinforcement learning (RL) into BFR. RL has demonstrated remarkable success in expanding performance boundaries across various domains, particularly in language models (Shao et al., 2024; Yu et al., 2025) and vision models (Fan et al., 2023; Liu et al., 2025; Wang et al., 2025; Yuan et al., 2025), by enabling diverse exploration strategies rather than deterministic outputs. Building on recent advances that successfully integrate RL with diffusion models (Liu et al., 2025; Xue et al., 2025)–where the denoising process is formulated as a Markov decision process–we leverage RL’s exploration capabilities alongside diffusion models’ inherent randomness to systematically search BFR’s solution space for optimal restorations that enhance both fidelity and perceptual quality.
@@
To this end, we introduce the first online RL framework to BFR: Likelihood-Regularized Policy Optimization (LRPO). Our framework utilizes a policy network to generate multiple diverse HQ candidate faces from each LQ input, effectively exploring the solution space rather than following a single deterministic trajectory. As shown in Figure 1 (left), this multi-candidate sampling strategy allows systematic exploration of potential solutions. A reward function evaluates each candidate, and relative advantages computed among candidates from the same input guide the policy network optimization. Moreover, we introduce three key innovations to enable effective RL optimization: (1) We design a composite reward function that evaluates restoration quality by incorporating human preferences, perceptual quality, and fidelity metrics. (2) To prevent reward hacking (Skalse et al., 2022; Amodei et al., 2016)–where the policy exploits reward signals while deviating from authentic facial distributions–we implement ground-truth (GT) guided likelihood regularization that anchors the policy to the true data manifold. (3) We propose a noise-level advantage assignment mechanism that weights the advantages according to the importance of each denoising step, ensuring more effective policy updates.
@@
Diffusion-based Blind Face Restoration. Blind Face Restoration (BFR) aims to recover high-quality face images from these degraded inputs while preserving identity consistency and perceptual quality. Recently, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have gained popularity due to their generative diversity and training stability. DR2 (Wang et al., 2023b) uses a diffusion model for degradation removal, followed by refinement through an enhancement module. DifFace (Yue & Loy, 2024) constructs a posterior distribution to map LQ images to HQ counterparts, utilizing the error-shrinkage property of pre-trained diffusion models for robust restoration. To accelerate training, LDM (Rombach et al., 2022) recommends training diffusion in the latent space. DiffBIR (Lin et al., 2024), built on LDM, employs ControlNet to guide restoration using low-quality faces as control signals. Some variations of diffusion models have been used for face restoration. InterLCM (Li et al., 2025) leverages the latent consistency model (LCM) to improve semantic consistency, restore images efficiently. FlowIE (Zhu et al., 2024) uses conditional rectified flow for faster inference with comparable restoration quality. However, diffusion models often suffer from poor identity consistency and loss of facial details in restored images. We find that intrinsic randomness of diffusion models can be leveraged through reinforcement learning with composite reward mechanisms to generate higher-quality, more detailed facial restorations.
RL in Vision Generation. Reinforcement learning has recently achieved remarkable success in improving large language model reasoning, particularly through policy gradient approaches such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024). In the text-to-image (T2I) generation field, many methods have explored incorporating policy gradient approaches (e.g., PPO) to align with human preferences. These methods explicitly cast diffusion denoising as a multi-step decision process and update the policy accordingly. DDPO (Black et al., 2023) improves alignment and aesthetics by optimizing rewards tied to human feedback. DPOK (Fan et al., 2023) studies online RL with Kullback–Leibler (KL) regularization on SD (Rombach et al., 2022). As a complementary approach to RL, Diffusion-DPO (Wallace et al., 2024) successfully adapts direct preference optimization to diffusion likelihoods, achieving significant improvements in human-preference alignment on SDXL (Podell et al., 2023). Beyond standard diffusion models, Flow-GRPO (Liu et al., 2025) is the first to bring GRPO into flow-matching model. Building on this, TempFlow-GRPO (He et al., 2025) further improves efficiency and stability by introducing trajectory branching and enabling process-level rewards without an intermediate reward model. Motivated by these successes, we present the first integration of policy gradient-based online RL into the BFR domain.
@@
BFR Problem Modeling. BFR is an ill-posed inverse problem. From a mathematical perspective, given a low-quality observation $\bm{c}_{\text{LQ}}$, the posterior distribution of its corresponding high-quality $\bm{x}_{0}$, denoted as $p(\bm{x}_{0}|\bm{c}_{\text{LQ}})$, has multiple feasible solutions (Menon et al., 2020). This posterior can be modeled as a mixture distribution:
$p(\bm{x}_{0}|\bm{c}_{\text{LQ}})=\sum_{k=1}^{K}w_{k}\cdot p_{k}(\bm{x}_{0}|\bm{c}_{\text{LQ}})$ (1)
Here, $K$ represents the number of possible high-quality solutions. Each distribution $p_{k}(\bm{x}_{0}|\bm{c}_{\text{LQ}})$ represents a real face distribution compatible with $\bm{c}_{LQ}$, having a specific identity, expression, or detail, with a peak at $\bm{\mu}_{k}$. The term $w_{k}$ denotes the probability of the $k$-th solution, where $\sum w_{k}=1$. This multi-solution nature poses a challenge for existing methods. Without dedicated exploration mechanisms, restoration models tend to converge toward average solutions, resulting in characteristically blurry faces that lack rich textural details (Lugmayr et al., 2020).
@@
The denoising process recovers data via the conditional distribution $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t},\bm{c})$, where $\bm{c}$ is a condition. In the DDIM (Song et al., 2020) framework, the network predicts the noise $\epsilon_{\theta}(\bm{x}_{t},t,\bm{c})$. The one-step denoising formula (from $t$ to t1t-1) is given by:
$\mu_{\theta}(\bm{x}_{t},t,\bm{c})$ $=\sqrt{\alpha_{t-1}}\cdot\frac{\bm{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon_{\theta}(\bm{x}_{t},t,\bm{c})}{\sqrt{\alpha_{t}}}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}(\bm{x}_{t},t,\bm{c}),$ (2)
$\quad\quad\quad\bm{x}_{t-1}=\mu_{\theta}(\bm{x}_{t},t,\bm{c})+\sigma_{t}\cdot\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(\bm{0},\mathbf{I}),$ (3)
@@
Fan et al. (2023) proposes that DDIM can be formulated as a Markov Decision Process (MDP) defined by the tuple $(S,A,\rho_{0},\pi,P,R)$. At time step $t$, the state is $\bm{s}_{t}\triangleq(\bm{c},t,\bm{x}_{t})$. The action is the denoised sample predicted by the model, $\bm{a}_{t}\triangleq\bm{x}_{t-1}$, with the policy defined as $\pi(\bm{a}_{t}|\bm{s}_{t})\triangleq\pi_{\theta}(\bm{x}_{t-1}|\bm{x}_{t},\bm{c})$. The state transition is deterministic, given by $P(\bm{s}_{t+1}|\bm{s}_{t},\bm{a}_{t})\triangleq(\delta_{\bm{c}},\delta_{\bm{a}_{t}})$, where $\delta_{\bm{c}}$ denotes the Dirac distribution at $\bm{c}$. The initial state distribution is $P_{0}(\bm{s}_{0})\triangleq(p(\bm{c}),\mathcal{N}(0,\mathbf{I}))$. The reward is provided only at the final step: $R(\bm{s}_{t},\bm{a}_{t})\triangleq r(\bm{x}_{0},\bm{c})$. When $\eta>0$, the DDIM supporting MDP can achieve reinforcement learning training.
@@ 4 Methodology
LRPO is an approach that enhances the BFR task using online RL, as shown in Figure 2. First, we initialize its policy network with an off-the-shelf diffusion-based face restoration model. Specifically, we employ DiffBIR (Lin et al., 2024) (our base model), a ControlNet-based (Zhang et al., 2023) approach that uses the LQ input as a control signal to guide restoration. Based on the GRPO (Shao et al., 2024) algorithm, we propose three core innovations for the BFR task: (1) We design a composite reward function that provides rewards for the diffusion denoising process. (2) We propose a ground-truth guided likelihood regularization term to penalize policy updates that deviate from real face data. (3) We develop a noise-aware advantage assignment mechanism to appropriately weight advantages based on denoising step significance.
@@ 4.1 Group Relative Policy Optimization
The optimization goal of reinforcement learning is to maximize the expected cumulative reward. Unlike PPO (Schulman et al., 2017), GRPO (Shao et al., 2024) samples a group of $G$ trajectories $\{\{\bm{x}_{T}^{(i)},\bm{x}_{T-1}^{(i)},\dots,\bm{x}_{0}^{(i)}\}\}_{i=1}^{G},$ from the policy $\pi_{\theta_{\text{old}}}$ , obtaining $G$ candidate reconstructions $\hat{\bm{x}}_{0}^{(i)}$, which are decoded from $\bm{x}_{0}^{(i)}$ in latent diffusion, and their rewards $r^{(i)}=r(\hat{\bm{x}}_{0}^{(i)},\bm{x}_{\text{GT}})$, where $\bm{x}_{\text{GT}}$ denotes the ground truth corresponding to $\hat{\bm{x}}_{0}^{(i)}$. Then, the advantage of the $i$-th sampled trajectory at time $t$ is calculated by normalizing the group-level rewards:
@@
$\mathcal{J}_{\text{DDIM-GRPO}}(\theta)=\mathbb{E}_{\bm{c}_{\text{LQ}}\sim p(\bm{c}_{\text{LQ}}),\{\hat{\bm{x}}^{(i)}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|\bm{c}_{\text{LQ}})}$ (5)
$\quad\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\Bigg[\min\left(r_{t}^{(i)}(\theta)\hat{A}_{t}^{(i)},\text{clip}\left(r_{t}^{(i)}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{t}^{(i)}\right)-\beta D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})\Bigg],$
@@ 4.2 Likelihood-Regularized Policy Optimization
Composite Reward Function. To effectively guide the optimization of the policy network $\pi_{\theta}$ and enable it to find a better balance in the complex Perception-Distortion Tradeoff (Blau & Michaeli, 2018), we design a multi-objective composite reward function, $R(\hat{\bm{x}}_{0},\bm{x}_{\text{GT}})$. This composite function measures the quality of generated face images $\hat{\bm{x}}_{0}$ from three complementary perspectives:
  • Human Preference Reward ($r_{\text{pref}}$): We employ a Face Reward Model (Wu et al., 2025), pre-trained on a human preference dataset, to score the overall realism and naturalness of facial details, ensuring alignment with human aesthetic preferences.
  • Perceptual Quality ($r_{\text{aq}}$): We incorporate CLIP-IQA (Wang et al., 2023a), a no-reference image quality assessment metric, to objectively measure perceptual quality using knowledge from pre-trained CLIP (Radford et al., 2021) models.
@@
$R(\hat{\bm{x}}_{0},\bm{x}_{\text{GT}})=\lambda_{1}r_{\text{pref}}(\hat{\bm{x}}_{0})+\lambda_{2}r_{\text{aq}}(\hat{\bm{x}}_{0})+\lambda_{3}r_{\text{fid}}(\hat{\bm{x}}_{0},\bm{x}_{\text{GT}})$ (7)
@@
$\mathcal{R}_{\text{likelihood}}=-\log{\pi}_{\theta}(\bm{x}_{t-1}^{\mathrm{GT}}|\bm{x}_{t},\bm{c}_{\text{LQ}})\propto\frac{\|\bm{x}_{t-1}^{\mathrm{GT}}-\mu_{\theta}(\bm{x}_{t},t,\bm{c}_{\text{LQ}})\|_{2}^{2}}{2\sigma_{t}^{2}}.$ (8)
@@
Noise-Level Advantage Assignment. Conventional GRPO-based approaches (Liu et al., 2025; Fan et al., 2023) treat all timesteps equally when assigning advantage weights, ignoring the inherently non-uniform importance of different steps in the diffusion process. To address this, inspired by previous work (He et al., 2025), we introduce a noise-level-aware advantage assignment approach that correlates advantage weights with the exploration magnitude achieved at each denoising step. In DDIM sampling, the single-step exploration radius from $\bm{x}_{t}$ to $\bm{x}_{t-1}$ is determined by the standard deviation $\sigma_{t}$ of the added noise. Therefore, we set the timestep weight $w_{t}$ proportional to $\sigma_{t}$:
$w_{t}\propto\sigma_{t},\quad\text{s.t.}\quad\frac{1}{T}\sum_{t=0}^{T-1}w_{t}=1$ (9)
@@
$\mathcal{J}_{\text{LRPO}}(\theta)=\mathbb{E}_{\bm{c}_{\text{LQ}}\sim p(\bm{c}_{\text{LQ}}),\{\hat{\bm{x}}^{(i)}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|\bm{c}_{\text{LQ}})}$ (11)
$\quad\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\Bigg[\min\left(r_{t}^{(i)}(\theta)\hat{A}_{t}^{(i)},\text{clip}\left(r_{t}^{(i)}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{t}^{(i)}\right)+\alpha\mathcal{R}_{\text{likelihood}}^{(i)}\Bigg],$
Compared to Eq. LABEL:eq:flow-grpo, we replace the KL divergence component with GT-guided likelihood regularization in our optimization objective.

### v2

# Enhancing Blind Face Restoration through Online Reinforcement Learning

@@ 1 Introduction
Blind Face Restoration (BFR), which aims to reconstruct high-quality (HQ) faces from low-quality (LQ) inputs with unknown degradations, has made rapid progress in recent years. Modern BFR methods typically exploit various types of priors to establish direct mappings from LQ to HQ. These priors can be categorized into several types: (1) geometric priors (e.g., facial landmarks [7, 18], parsing maps [5], and component heatmaps [46]) that provide structural guidance; (2) generative priors [36, 4, 44] derived from pre-trained models such as StyleGAN [14, 15], which enable realistic detail reconstruction; (3) discrete codebook priors [9, 52] improve restoration fidelity; and (4) diffusion priors [41, 21, 6, 48, 39] is built upon the denoising and noise-adding diffusion model. Diffusion models offer distinct advantages including robust generative capability, stable optimization, and superior control over output diversity, making them particularly effective for producing high-quality, visually pleasing face restorations.
However, despite the advantages of diffusion priors, BFR remains fundamentally challenging. The task is an ill-posed inverse problem where a single LQ input can be corresponded to multiple plausible HQ solutions, making it difficult to determine the optimal restoration. Current methods are constrained by their deterministic nature. They learn a fixed one-to-one mapping that produces a single output without considering alternative solutions. This lack of exploration within the vast solution space prevents these methods from discovering potentially superior restorations, leading to suboptimal results [51].
To address these exploration limitations, we propose incorporating reinforcement learning (RL) into BFR. RL has demonstrated remarkable success across various domains, particularly in language models [29, 45] and vision models [8, 22, 35, 47], by enabling diverse exploration strategies rather than deterministic outputs. Building on recent advances that successfully integrate RL with diffusion models [22, 42]—where the denoising process is formulated as a Markov decision process—we leverage RL’s exploration capabilities alongside diffusion models’ inherent randomness to systematically search BFR’s solution space for optimal restorations that enhance both fidelity and perceptual quality.
@@
To this end, we introduce the first online RL framework to BFR: Likelihood-Regularized Policy Optimization (LRPO). Our framework utilizes a policy network to generate multiple diverse HQ candidate faces from each LQ input, effectively exploring the solution space rather than following a single deterministic trajectory. As shown in Figure 1 (left), this multi-candidate sampling strategy allows systematic exploration of potential solutions. A reward function evaluates each candidate, and relative advantages computed among candidates from the same input guide the policy network optimization. Moreover, we introduce three key innovations to enable effective RL optimization: (1) We design a composite reward function that evaluates restoration quality by incorporating human preferences, perceptual quality, and fidelity metrics. (2) To prevent reward hacking [30, 1]—where the policy exploits reward signals while deviating from authentic facial distributions–we implement ground-truth (GT) guided likelihood regularization that anchors the policy to the true data manifold. (3) We propose a noise-level advantage assignment mechanism that weights the advantages according to the importance of each denoising step, ensuring more effective policy updates.
@@
Diffusion-based Blind Face Restoration. Blind Face Restoration (BFR) aims to recover high-quality face images from these degraded inputs while preserving identity consistency and perceptual quality. Recently, diffusion models [31, 12] have gained popularity due to their generative diversity and training stability. DR2 [39] uses a diffusion model for degradation removal, followed by refinement through an enhancement module. DifFace [48] constructs a posterior distribution to map LQ images to HQ counterparts, utilizing the error-shrinkage property of pre-trained diffusion models for robust restoration. To accelerate training, LDM [27] recommends training diffusion in the latent space. DiffBIR [21], built on LDM, employs ControlNet to guide restoration using low-quality faces as control signals. Some variations of diffusion models have been used for face restoration. InterLCM [20] leverages the latent consistency model (LCM) to improve semantic consistency, restore images efficiently. FlowIE [53] uses conditional rectified flow for faster inference with comparable restoration quality. However, diffusion models often suffer from poor identity consistency and loss of facial details in restored images. We find that randomness of diffusion models can be leveraged through reinforcement learning with composite reward mechanisms to generate higher-quality facial restorations.
RL in Vision Generation. Reinforcement learning has recently achieved remarkable success in improving large language model reasoning, particularly through policy gradient approaches such as PPO [28] and GRPO [29]. In the text-to-image (T2I) generation field, many methods have explored incorporating policy gradient approaches (e.g., PPO) to align with human preferences. These methods explicitly cast diffusion denoising as a multi-step decision process and update the policy accordingly. DDPO [2] improves alignment and aesthetics by optimizing rewards tied to human feedback. DPOK [8] studies online RL with Kullback–Leibler (KL) regularization on SD [27]. As a complementary approach to RL, Diffusion-DPO [33] successfully adapts direct preference optimization to diffusion likelihoods, achieving significant improvements in human-preference alignment on SDXL [25]. Beyond standard diffusion models, Flow-GRPO [22] is the first to bring GRPO into flow-matching model. Building on this, TempFlow-GRPO [10] further improves efficiency and stability by introducing trajectory branching and enabling process-level rewards without an intermediate reward model. Motivated by these successes, we present the first integration of policy gradient-based online RL into the BFR domain.
@@
BFR Problem Modeling. BFR is an ill-posed inverse problem. From a mathematical perspective, given a low-quality observation $\bm{c}_{\text{LQ}}$, the posterior distribution of its corresponding high-quality $\bm{x}_{0}$, denoted as $p(\bm{x}_{0}|\bm{c}_{\text{LQ}})$, has multiple feasible solutions [24]. This posterior can be modeled as a mixture distribution:
$p(\bm{x}_{0}|\bm{c}_{\text{LQ}})=\sum_{k=1}^{K}w_{k}\cdot p_{k}(\bm{x}_{0}|\bm{c}_{\text{LQ}}).$ (1)
Here, $K$ represents the number of possible high-quality solutions. Each distribution $p_{k}(\bm{x}_{0}|\bm{c}_{\text{LQ}})$ represents a real face distribution compatible with $\bm{c}_{LQ}$, having a specific identity, expression, or detail, with a peak at $\bm{\mu}_{k}$. The $w_{k}$ denotes the probability of the $k$-th solution, where $\sum w_{k}=1$. This multi-solution nature poses a challenge for existing methods. Without dedicated exploration mechanisms, restoration models tend to converge toward average solutions, resulting in blurry faces with weakened textural detail [23].
@@
The denoising process recovers data via the conditional distribution $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t},\bm{c})$, where $\bm{c}$ is a condition. In the DDIM [32] framework, the network predicts the noise $\epsilon_{\theta}(\bm{x}_{t},t,\bm{c})$. The one-step denoising formula (from $t$ to t1t-1) is given by:
$\begin{split}\mu_{\theta}(\bm{x}_{t},t,\bm{c})&=\sqrt{\alpha_{t-1}}\cdot\frac{\bm{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon_{\theta}(\bm{x}_{t},t,\bm{c})}{\sqrt{\alpha_{t}}}\\ &\quad+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}(\bm{x}_{t},t,\bm{c}),\end{split}$ (2)
$\bm{x}_{t-1}$ $=\mu_{\theta}(\bm{x}_{t},t,\bm{c})+\sigma_{t}\cdot\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(\bm{0},\mathbf{I}).$ (3)
@@
[8] proposes that DDIM can be formulated as a Markov Decision Process (MDP) defined by the tuple $(S,A,\rho_{0},\pi,P,R)$. At time step $t$, the state is $\bm{s}_{t}\triangleq(\bm{c},t,\bm{x}_{t})$. The action is the denoised sample predicted by the model, $\bm{a}_{t}\triangleq\bm{x}_{t-1}$, with the policy defined as $\pi(\bm{a}_{t}|\bm{s}_{t})\triangleq\pi_{\theta}(\bm{x}_{t-1}|\bm{x}_{t},\bm{c})$. The state transition is deterministic, given by $P(\bm{s}_{t+1}|\bm{s}_{t},\bm{a}_{t})\triangleq(\delta_{\bm{c}},\delta_{\bm{a}_{t}})$, where $\delta_{\bm{c}}$ denotes the Dirac distribution at $\bm{c}$. The initial state distribution is $P_{0}(\bm{s}_{0})\triangleq(p(\bm{c}),\mathcal{N}(0,\mathbf{I}))$. The reward is provided only at the final step: $R(\bm{s}_{t},\bm{a}_{t})\triangleq r(\bm{x}_{0},\bm{c})$. When $\eta>0$, the DDIM MDP can achieve reinforcement learning training.
@@ 4 Methodology
LRPO is an approach that enhances the BFR task using online RL, as shown in Figure 2. First, we initialize its policy network with an off-the-shelf diffusion-based face restoration model. Specifically, we employ DiffBIR [21] (our base model), a ControlNet-based [49] approach that uses the LQ input as a control signal to guide restoration. Based on the GRPO [29] algorithm, we propose three core innovations for the BFR task: (1) We design a composite reward function that provides rewards for the diffusion denoising process. (2) We propose a ground-truth guided likelihood regularization term to penalize policy updates that deviate from real face data. (3) We develop a noise-aware advantage assignment mechanism to appropriately weight advantages based on denoising step significance. LRPO is designed to post-train off-the-shelf face restoration models via RL, thereby extending their performance frontier. After training, the RL component is discarded, so no additional inference overhead is introduced to the base model.
@@ 4.1 Group Relative Policy Optimization
The optimization goal of reinforcement learning is to maximize the expected cumulative reward. Unlike PPO [28], GRPO [29] samples a group of $G$ trajectories $\{\{\bm{x}_{T}^{(i)},\bm{x}_{T-1}^{(i)},\dots,\bm{x}_{0}^{(i)}\}\}_{i=1}^{G},$ from the policy $\pi_{\theta_{\text{old}}}$ , obtaining $G$ candidate reconstructions $\hat{\bm{x}}_{0}^{(i)}$, which are decoded from $\bm{x}_{0}^{(i)}$ in latent diffusion, and their rewards $r^{(i)}=r(\hat{\bm{x}}_{0}^{(i)},\bm{x}_{\text{GT}})$, where $\bm{x}_{\text{GT}}$ denotes the ground truth corresponding to $\hat{\bm{x}}_{0}^{(i)}$. Then, the advantage of the $i$-th sampled trajectory at time $t$ is calculated by normalizing the group-level rewards:
@@
$\mathcal{J}_{\text{DDIM-GRPO}}(\theta)=\mathbb{E}_{\bm{c}_{\text{LQ}},\{\hat{\bm{x}}^{(i)}\}_{i=1}^{G}}\left[\mathcal{L}(\theta)\right],$ (5)
$\mathcal{L}(\theta)$ $=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\Bigg[\min\left(r_{t}^{(i)}(\theta)\hat{A}_{t}^{(i)},\right$. (6)
$\left.\text{clip}\left(r_{t}^{(i)}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{t}^{(i)}\right)-\beta D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})\Bigg].$
@@ 4.2 Likelihood-Regularized Policy Optimization
Composite Reward Function. To effectively guide the optimization of the policy network $\pi_{\theta}$ and enable it to find a better balance in the complex Perception-Distortion Tradeoff [3], we design a multi-objective composite reward function, $R(\hat{\bm{x}}_{0},\bm{x}_{\text{GT}})$. This composite function measures the quality of generated face images $\hat{\bm{x}}_{0}$ from three complementary perspectives:
  • Human Preference Reward ($r_{\text{pref}}$): We employ a Face Reward Model [40], pre-trained on a human preference dataset, to score the overall realism and naturalness of facial details, ensuring alignment with human aesthetic preferences.
  • Perceptual Quality ($r_{\text{aq}}$): We incorporate CLIP-IQA [34], a no-reference image quality assessment metric, to objectively measure perceptual quality using knowledge from pre-trained CLIP [26] models.
@@
$R(\hat{\bm{x}}_{0},\bm{x}_{\text{GT}})=\lambda_{1}r_{\text{pref}}(\hat{\bm{x}}_{0})+\lambda_{2}r_{\text{aq}}(\hat{\bm{x}}_{0})+\lambda_{3}r_{\text{fid}}(\hat{\bm{x}}_{0},\bm{x}_{\text{GT}}).$ (8)
@@
$\mathcal{R}_{\text{likelihood}}=-\log{\pi}_{\theta}(\bm{x}_{t-1}^{\mathrm{GT}}|\bm{x}_{t},\bm{c}_{\text{LQ}}),$ (9)
$\mathcal{R}_{\text{likelihood}}\propto\frac{\|\bm{x}_{t-1}^{\mathrm{GT}}-\mu_{\theta}(\bm{x}_{t},t,\bm{c}_{\text{LQ}})\|_{2}^{2}}{2\sigma_{t}^{2}}.$ (10)
@@
Noise-Level Advantage Assignment. Conventional GRPO-based approaches [22, 8] treat all timesteps equally when assigning advantage weights, ignoring the inherently non-uniform importance of different steps in the diffusion process. To address this, inspired by previous work [10], we introduce a noise-level-aware advantage assignment approach that correlates advantage weights with the exploration magnitude achieved at each denoising step. In DDIM sampling, the single-step exploration radius from $\bm{x}_{t}$ to $\bm{x}_{t-1}$ is determined by the standard deviation $\sigma_{t}$ of the added noise. Therefore, we set the timestep weight $w_{t}$ proportional to $\sigma_{t}$:
$w_{t}\propto\sigma_{t},\quad\text{s.t.}\quad\frac{1}{T}\sum_{t=0}^{T-1}w_{t}=1.$ (11)
@@
$\mathcal{J}_{\text{LRPO}}(\theta)=\mathbb{E}_{\bm{c}_{\text{LQ}},\{\hat{\bm{x}}^{(i)}\}_{i=1}^{G}}\left[\mathcal{L}_{\text{LRPO}}(\theta)\right],$ (13)
$\mathcal{L}$ ${}_{\text{LRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\Bigg[\min\left(r_{t}^{(i)}(\theta)\hat{A}_{t}^{(i)},\right$. (14)
$\left.\text{clip}\left(r_{t}^{(i)}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{t}^{(i)}\right)+\alpha\mathcal{R}_{\text{likelihood}}^{(i)}\Bigg].$
Compared to Eq. 5, we replace the KL divergence component with GT-guided likelihood regularization in our optimization objective.